the area of a carpet is 308 sq m .If its length is 2 meters less and its width 6 meters more,the carpet will be square.Find the dimensions of the carpet.
P.S Solving Quadratic Equations by Completing the Square is the topic.

Answer by amoresroy(361)

the area of a carpet is 308 sq m .If its length is 2 meters less and its width 6 meters more,the carpet will be square.Find the dimensions of the carpet.
P.S Solving Quadratic Equations by Completing the Square is the topic.
Let x = the width of the carpet
y = the length of the carpet
1st equation
xy = 308
y = 308/x
2nd equation
y-2 = x+6
since y=308/x in 1st equation, the 2nd equation becomes
308/x - 2 = x+6
combine like terms
308/x -x - 8 = 0
multiply both sides by x
308 - x^2 -8x = 0
multiply both sides by -1 and rearrange
x^2 + 8x - 308 = 0
(x+22)(x-14) = 0
x-14 = 0
x = 14
So y = 300/14 = 22
The width of the carpet is 14 meters and the length of the carpet is 22 meters.